Characteristics of Production Process, Cost Measurement

Vince Melders, of EcoScape Company, designs and installs custom lawn and garden irrigation systems for homes and businesses throughout the state. Each job is different, requiring different materials and labor for installing the systems. EcoScape estimated the following for the year:

Number of direct labor hours 6,600
Direct labor cost $85,800
Overhead cost $52,800

During the year, the following actual amounts were experienced:

Number of direct labor hours 6,145
Direct labor incurred $86,030
Overhead incurred $53,817.91

Required:

1. Should EcoScape use process costing or job-order costing?

2. If EcoScape uses a normal costing system and overhead is applied on the basis of direct labor hours,

What is the overhead rate? Round your answer to the nearest cent.

$ per direct labor hour

What is the average actual wage rate?

$ per direct labor hour

What is the cost of an installation that takes $3,270 of direct materials and 20 direct labor hours?

$
Answer #1

1) EcoScape use Job order costing

2a) Overhead rate = Estimated overhead/Estimated labor hour = 52800/6600 = 8 per labor hour

2b) Average actual wages rate = 86030/6145 = 14 per labor hour

2c) Cost on installation

Direct material 3270
Direct labor (20*14) 280
Overhead applied (20*8) 160
Total cost of installation 3710